A uniform disc having a mass of 5kg and radius of gyration of 150mm is mounted on one end of a horizontal arm of length 200mm. the other end of the arm can rotate freely in a bearing. The disc is given a, clockwise spin of 250 rpm when seen from the end. Determine the motion of disc if the arm remains horizontal. [5] 


1
Expert's answer
2021-10-05T13:47:56-0400

"m=5kg, l=200mm=0.2m, k=150mm=0.15m, N=250rpm"

"I=mk^2=5\times 0.15^2=0.1125kg.m^2"

"\omega=\frac{2\pi \times 250}{60}=26.18rad/s"

"C=I\omega \omega_p"

"mgl=I\omega \omega_p"

"5\times 9.18\times 0.2=0.1125\times 26.18\times \omega_p"

"\omega_p=3.33rad/s"